OpenStudy (anonymous):

Fan and medal? Explain how to solve the following system of equations. What is the solution to the system? 2x + 2y + z = -5 3x + 4y + 2z = 0 x + 3y + 2z = 1

OpenStudy (anonymous):

ok even though z doesn't have a number in front of it in math we always put a 1 in front of it so it would be like this: 1z just to help you lets put that process to all of them: to make life easier: 2x + 2y + 1z = -5 3x + 4y + 2z = 0 1x + 3y + 2z = 1 now just solve

OpenStudy (anonymous):

6x + 9y + 5z = -4

OpenStudy (anonymous):

now explain how you got that answer

OpenStudy (anonymous):

Okay, thank you :) very much

OpenStudy (anonymous):

Your very welcome👏

OpenStudy (anonymous):

Ahaha ^.^ thank you!